OpenStudy (xoomd):

PLEASE HELP, WILL FAN AND MEDAL Choose the system of equations which matches the following graph: A: x + 2y = −10 3x − y = 12 B: x − 2y = −10 3x − y = 12 C: x + 2y = −10 3x + y = 12 D: x − 2y = −10 3x + y = 12

OpenStudy (xoomd):

ill post a pic of the graph

OpenStudy (ajspeller):

Find the coordinates of the location where the graphs intersect. Then substitute in the values of x and y into each of the equations.

OpenStudy (xoomd):

So, I got the cordinates 2,-6 do you agree with those?

OpenStudy (ajspeller):

Yes

OpenStudy (xoomd):

im confused by what i do with those points now

OpenStudy (ajspeller):

Now substitute in 2 for x and -6 for y in each the equations. x+2y=10 (2) + 2(-6) = 10 Evaluate each pair of equations to see if both statement are true. That would be the correct response.

OpenStudy (xoomd):

Is it B?

OpenStudy (xoomd):

nevermind it is A